# Personalized Pricing Through Strategic User Profiling in Social Networks

### v1

@@ VIII-B Partially Connected Social Networks
Figure 11: A subnetwork of the Facebook social network sourced from [37], and an illustration of different users’ social activity levels in the PBE (indicated by the equilibrium valuation thresholds $v_{i}^{*}$ for each user $i\in\mathcal{N}$) with a profiling accuracy $\delta=0.7$. Here, we consider that the valuation of each subnetwork user follows a uniform distribution over $[0,40]$.
How should users manage online activities when interacting within a partially connected social network? Moreover, considering the diverse and distinctive network positions among users, can the degree of each user node alone suffice to indicate their actions? To answer these questions, we will first present key structural results for the PBE of this new extension in Appendix P, based on which we then conduct empirical experiments utilizing Facebook social network data from [37].
For our empirical studies, we utilize Facebook social network data from [37] and focus on a typical induced subnetwork of the original Facebook network in [37]. Specifically, the subnetwork consists of $100$ user nodes and $215$ edges, with an average degree of $4.3$. Fig. 11(a) provides a visualization of this subnetwork. Furthermore, in Fig. 11(b), we graphically illustrate users’ social activity levels in the PBE, as indicated by their equilibrium valuation thresholds $\{v^{*}_{i},\forall i\in\mathcal{N}\}$, with a profiling accuracy $\delta=0.7$. In particular, the users who appear to be more connected and central in the network have higher valuation thresholds and are more likely to be active online.
@@ IX Robustness Beyond Uniform Valuation Distribution
Assumption 1 presents a set of sufficient conditions to ensure the existence and uniqueness of PBE beyond the uniform distribution. Typical examples that fall under these conditions include the uniform distribution, the exponential distribution, and certain beta distributions. [^9] For instance, the CDF of beta distributions with shape parameters $\alpha=1$, $1<\beta<2$ or $1<\alpha<2$, $\beta=1$ is concave. Particularly, as in the literature on price discrimination (e.g., [16, 38]), we also assume the concavity of the benchmark revenue function $p_{0}(1-F(p_{0}))$. This indicates that the seller’s marginal revenue without user profiling decreases in the uniform price. This condition ensures the existence of a unique global maximum in the no-profiling benchmark case, which enables tractable analysis through the first-order condition. We hereafter denote the unique solution to $\arg\max_{p_{0}}p_{0}(1-F(p_{0}))$ as $\hat{p}_{0}$, i.e.,
@@
We are also interested in whether such non-monotonicity in users’ equilibrium social behaviors carries through to more general valuation distributions. Below, we conduct an empirical study based on real-world consumer profiles sourced from an annually-updated consumer dataset in [39]. Specifically, our analysis focuses on handbag purchases and the corresponding $92$ consumer profiles. To start, we utilize these empirical data to estimate the users’ valuation distribution, which follows a normal distribution $\mathcal{N}(57.84,20.25^{2})$ truncated over $[20,100]$ (see more details in Appendix Q-A). Based on this estimated market information, we proceed to derive the equilibrium behaviors of both (sampled) users and the seller in Fig. 13. Please refer to Appendix P for a complete version of this empirical study, where we delve into the PBE outcome in more detail and further examine the seller’s empirical revenue gain.
@@
Figure 13: The fraction of socially active users $F(v^{\ast})$ and the seller’s uniform price $p_{0}^{*}$ in the PBE (whenever it exists) based on real-world dataset [39].
@@ C-A Step I: Game Formulation and Solution Existence
Next, we prove that the social interaction game defined above is a supermodular game [32, 40].
@@
Any Bayesian game is itself a supermodular game when its underlying certainty game (for any valuation realization) is supermodular (see Theorem 6.1 in [41]); thus we will next show the supermodularity of the underlying certainty game for $\Gamma$ to complete the proof. According to the definition of the supermodular game (see Theorem 4 in [32]), proofs can be done by verifying that (i) the action set $\mathcal{X}_{i}\triangleq[0,1]\in\mathbb{R}$ for each user $i$ is a compact subset of real number and thus a complete lattice, (ii) the utility function $\tilde{\pi}_{i}(x_{i},\boldsymbol{x}_{-i})$ is supermodular in $x_{i}$ (for fixed $\boldsymbol{x}_{-i}$) since $x_{i}$ is a single variable instead of multidimensional vector, and (iii) $\tilde{\pi}_{i}(x_{i},\boldsymbol{x}_{-i})$ is twice continuously differentiable in $x_{i}$ over $[0,1]$ and has increasing differences in $x_{i}$ and $x_{j}$, $\forall j\neq i$, i.e.,
@@
This follows from Proposition A1 based on the properties of the supermodular game (see Theorem 5 in [32]). ∎
@@ Appendix E Proof for Proposition 2
$=\sum_{m=0}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})mF(v^{\dagger})^{m-1}\left(1-F(v^{\dagger})\right)^{n-1-m}f(v^{\dagger})$
$\quad-\sum_{m=0}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})(n-1-m)F(v^{\dagger})^{m}\left(1-F(v^{\dagger})\right)^{n-2-m}f(v^{\dagger}),$
$=\sum_{m=1}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})mF(v^{\dagger})^{m-1}\left(1-F(v^{\dagger})\right)^{n-1-m}f(v^{\dagger})$
$\quad-\sum_{m=0}^{n-2}\tbinom{n-1}{m}\ln(m+\omega_{0})(n-1-m)F(v^{\dagger})^{m}\left(1-F(v^{\dagger})\right)^{n-2-m}f(v^{\dagger}),$
@@
$\ln(m+\omega_{0})-2\ln(m-1+\omega_{0})+\ln(m-2+\omega_{0})$
$=$ $\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)-\left(\ln(m-1+\omega_{0})-\ln(m-2+\omega_{0})\right)<0;$
@@ G-B Step 2: Seller’s Uniform Pricing Analysis
To constitute the PBE, users’ social activity decisions should be consistent with the seller’s belief that is updated according to Bayes’ theorem (i.e., belief consistency [18]). Moreover, anticipating the seller’s uniform price $p_{0}$ in Stage II, users must have no incentive to deviate from the equilibrium strategies of $x_{i}^{*}=1,\forall i\in\mathcal{N}$. This fact then helps guide our proof in this subsection. Rewrite the final payoff for each user $i$ in (4) here:
@@ H-A Impact of $\delta$ on $v^{*}$
    and $\partial^{2}\Theta(v^{*},\delta)/\partial\delta^{2}=2v^{*}>0$, which suggests that $\partial\Theta(v^{*},\delta)/\partial\delta$ increases with $\delta$ on $\left(2\ln(n-1+\omega_{0})/\bar{v},1\right)$. We also calculate the following second-order partial derivative:
@@ I-B Proof for Proposition 5
$\sum_{m=1}^{n-1}\sum_{k=1}^{n-1}\frac{(n-2)!}{(n-1-m)!(m-1)!}\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)\leq\\ \sum_{m=1}^{n-1}\sum_{k=1}^{n-1}\frac{(n-2)!}{(n-1-k)!(k-1)!}\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)$ (A67)
@@
$\tilde{\Pi}_{0}\triangleq\mathbb{E}_{i\sim\mathcal{N}_{0}}\left(\sum_{i\in\mathcal{N}_{0}}p_{0}d_{i}\right)=|\mathcal{N}|\text{Prob}\left[i\in\mathcal{N}_{0}\right]\cdot\mathbb{E}_{v_{i}\sim f(\cdot|i\in\mathcal{N}_{0})}\left(p_{0}\mathbbm{1}(v_{i}\geq p_{0})\right)=|\mathcal{N}|\left(1-\delta\frac{v^{*}}{\bar{v}}\right)p_{0}\left(1-\frac{1-\delta}{\bar{v}-\delta v^{*}}p_{0}\right),$ (A76)
@@
$\left((\delta-1)\tilde{L}^{\prime}-(2-\delta)\tilde{J}\right)\frac{\bar{v}}{\delta}+4\frac{1-\delta}{\delta^{2}}\tilde{J}\tilde{L}^{\prime}\leq\left(\left((\delta-1)\tilde{L}^{\prime}-(2-\delta)\tilde{J}\right)\frac{\bar{v}}{\delta}+4\frac{1-\delta}{\delta^{2}}\tilde{J}\tilde{L}^{\prime}\right)|_{\bar{v}=2\frac{\tilde{J}}{\delta}}=\frac{2\tilde{J}\tilde{L}^{\prime}}{\delta^{2}}\left((1-\delta)-(2-\delta)\frac{\tilde{J}}{\tilde{L}^{\prime}}\right)<0$ (A93)
@@
In the extended four-stage model, the seller sets the uniform price $p_{0}^{e}$ to maximize the total expected revenue $\tilde{\Pi}$ in (A77), i.e., $p_{0}^{e}=\arg\max_{p_{0}\in[0,\bar{v}]}\tilde{\Pi}(p_{0},v^{*}(p_{0}))$. Also, notice that $p_{0}^{e}$ is uniquely determined (see Appendix K-B3). It then follows that $\tilde{\Pi}^{e}\geq\tilde{\Pi}^{*}$ since $p_{0}^{e}\geq p_{0}^{*}$. Hereafter, we use the superscript $e$ to refer to the extended four-stage model and * for the original three-stage model. Then we turn to the expected profiled revenue $\tilde{\Pi}_{1}$ in (A75), and we will have $\tilde{\Pi}_{1}^{e}\geq\tilde{\Pi}_{1}^{*}$ since $v^{e}\geq v^{*}$. For the expected non-profiled revenue $\tilde{\Pi}_{0}$ in (A76), we have $\tilde{\Pi}_{0}^{e}=\tilde{\Pi}_{0}^{*}$ due to $v^{e}=v^{*}$ and $p_{0}^{e}=p_{0}^{*}$ if $\bar{v}/2\leq 1/\delta\cdot\ln\left(n-1+\omega_{0}\right)$ (Case I of Proposition 8). Otherwise, we have $\partial^{3}\tilde{\Pi}_{0}(v)/\partial v^{3}=\partial^{3}\tilde{\Pi}(v)/\partial v^{3}<0$ according to Lemma A12. Thus, $\partial\tilde{\Pi}_{0}(v)/\partial v$ is concave in $v$ over $v^{o},\bar{v})$ and $\partial\tilde{\Pi}_{0}(v)/\partial v=\partial\tilde{\Pi}(v)/\partial v-\partial\tilde{\Pi}_{1}(v)/\partial v=\partial\tilde{\Pi}(v)/\partial v-|\mathcal{N}|\delta v/\bar{v}$. Moreover, we have that at $v=v^{o}$:
@@
In this appendix, we evaluate the performance of our proposed pricing mechanism for the seller, through which we will understand the power of harnessing social network benefits (see Section VI-B). To evaluate the performance, we introduce two additional benchmarks adapted from the existing literature (e.g., [16, 22, 35, 25]). Specifically, we outline the two benchmarks for the pricing mechanism along with the one we proposed below.
  • Profile-independent Pricing (PIP) (e.g., [16, 22, 35]): The first benchmark involves the seller’s optimal pricing without differentiation based on user profiles. Variations of this benchmark are commonly adopted in the literature on behavior-based price discrimination. In this case, the seller charges a uniform price of $\bar{v}/2$ to all users.
  • Traditional-profiling-based Pricing (TPP) (e.g., [22, 25]): Another benchmark is wherein the seller personalizes price offerings through traditional user profiling. In this case, users voluntarily disclose private information without considering any social network benefits. Detailed equilibrium analyses are provided in Proposition 7, also named no-social-network benchmark there.
@@
In this appendix, we apply our proposed pricing mechanism to real-world consumer profiles, utilizing an annually-updated consumer behavior and shopping habits dataset from [39]. Recall that most of our previous analyses have assumed a uniform distribution for user valuations. However, as we will demonstrate later, this empirical experiment using more general distributions further confirms the robustness of our previous pricing methods and major insights.
@@
Figure A3: The user valuation distribution, the fraction of socially active users and the seller’s optimal uniform price $p_{0}^{*}$ in the PBE (whenever it exists), and the seller’s empirical revenue based on a real-world dataset from [39].
@@
The dataset from [39] offers consumer information such as demographic details, purchased items, purchase amounts, and previous purchase history. Our analysis focuses specifically on handbag purchases without discounts applied and the corresponding $92$ consumer profiles. [^10] We filter the dataset using the criteria ”Item Purchased = Handbag” AND ”Discount Applied = No.” In particular, we focus on purchases made without any promotional discounts applied. This allows us to exclude the potential psychological effects of price discounts on consumer perceptions and purchase intentions [42, 43]. Moreover, we associate each consumer’s purchase amount with his product valuation, representing the monetary value of his transaction on the purchased item denoted in United States Dollars (USD). To incorporate the social network benefit in our modeling framework, we further hypothetically consider these $92$ consumers as social network users within a community.

### v3

@@ VIII-B Partially Connected Social Networks
Figure 11: A subnetwork of the Facebook social network sourced from [snapnets], and an illustration of different users’ social activity levels in the PBE (indicated by the equilibrium valuation thresholds $v_{i}^{*}$ for each user $i\in\mathcal{N}$) with a profiling accuracy $\delta=0.7$. Here, we consider that the valuation of each subnetwork user follows a uniform distribution over $[0,40]$.
How should users manage online activities when interacting within a partially connected social network? Moreover, considering the diverse and distinctive network positions among users, can the degree of each user node alone suffice to indicate their actions? To answer these questions, we will first present key structural results for the PBE of this new extension in Appendix P, based on which we then conduct empirical experiments utilizing Facebook social network data from [snapnets].
For our empirical studies, we utilize Facebook social network data from [snapnets] and focus on a typical induced subnetwork of the original Facebook network in [snapnets]. Specifically, the subnetwork consists of $100$ user nodes and $215$ edges, with an average degree of $4.3$. Fig. 11(a) provides a visualization of this subnetwork. Furthermore, in Fig. 11(b), we graphically illustrate users’ social activity levels in the PBE, as indicated by their equilibrium valuation thresholds $\{v^{*}_{i},\forall i\in\mathcal{N}\}$, with a profiling accuracy $\delta=0.7$. In particular, the users who appear to be more connected and central in the network have higher valuation thresholds and are more likely to be active online.
@@ IX Robustness Beyond Uniform Valuation Distribution
Assumption 1 presents a set of sufficient conditions to ensure the existence and uniqueness of PBE beyond the uniform distribution. Typical examples that fall under these conditions include the uniform distribution, the exponential distribution, and certain beta distributions. [^9] For instance, the CDF of beta distributions with shape parameters $\alpha=1$, $1<\beta<2$ or $1<\alpha<2$, $\beta=1$ is concave. Particularly, as in the literature on price discrimination (e.g., [conitzer2012hide, fudenberg2006behavior]), we also assume the concavity of the benchmark revenue function $p_{0}(1-F(p_{0}))$. This indicates that the seller’s marginal revenue without user profiling decreases in the uniform price. This condition ensures the existence of a unique global maximum in the no-profiling benchmark case, which enables tractable analysis through the first-order condition. We hereafter denote the unique solution to $\arg\max_{p_{0}}p_{0}(1-F(p_{0}))$ as $\hat{p}_{0}$, i.e.,
@@
We are also interested in whether such non-monotonicity in users’ equilibrium social behaviors carries through to more general valuation distributions. Below, we conduct an empirical study based on real-world consumer profiles sourced from an annually-updated consumer dataset in [consumer]. Specifically, our analysis focuses on handbag purchases and the corresponding $92$ consumer profiles. To start, we utilize these empirical data to estimate the users’ valuation distribution, which follows a normal distribution $\mathcal{N}(57.84,20.25^{2})$ truncated over $[20,100]$ (see more details in Appendix Q-A). Based on this estimated market information, we proceed to derive the equilibrium behaviors of both (sampled) users and the seller in Fig. 13. Please refer to Appendix P for a complete version of this empirical study, where we delve into the PBE outcome in more detail and further examine the seller’s empirical revenue gain.
@@
Figure 13: The fraction of socially active users $F(v^{\ast})$ and the seller’s uniform price $p_{0}^{*}$ in the PBE (whenever it exists) based on real-world dataset [consumer].
@@ C-A Step I: Game Formulation and Solution Existence
Next, we prove that the social interaction game defined above is a supermodular game [1990Rationalizability, topkis1998supermodularity].
@@
Any Bayesian game is itself a supermodular game when its underlying certainty game (for any valuation realization) is supermodular (see Theorem 6.1 in [vives1990nash]); thus we will next show the supermodularity of the underlying certainty game for $\Gamma$ to complete the proof. According to the definition of the supermodular game (see Theorem 4 in [1990Rationalizability]), proofs can be done by verifying that (i) the action set $\mathcal{X}_{i}\triangleq[0,1]\in\mathbb{R}$ for each user $i$ is a compact subset of real number and thus a complete lattice, (ii) the utility function $\tilde{\pi}_{i}(x_{i},\boldsymbol{x}_{-i})$ is supermodular in $x_{i}$ (for fixed $\boldsymbol{x}_{-i}$) since $x_{i}$ is a single variable instead of multidimensional vector, and (iii) $\tilde{\pi}_{i}(x_{i},\boldsymbol{x}_{-i})$ is twice continuously differentiable in $x_{i}$ over $[0,1]$ and has increasing differences in $x_{i}$ and $x_{j}$, $\forall j\neq i$, i.e.,
@@
This follows from Proposition A1 based on the properties of the supermodular game (see Theorem 5 in [1990Rationalizability]). ∎
@@ Appendix E Proof for Proposition 2
$=\sum_{m=0}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})mF(v^{\dagger})^{m-1}\left(1-F(v^{\dagger})\right)^{n-1-m}f(v^{\dagger})-\sum_{m=0}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})(n-1-m)F(v^{\dagger})^{m}\left(1-F(v^{\dagger})\right)^{n-2-m}f(v^{\dagger}),$
$=\sum_{m=1}^{n-1}\tbinom{n-1}{m}\ln(m+\omega_{0})mF(v^{\dagger})^{m-1}\left(1-F(v^{\dagger})\right)^{n-1-m}f(v^{\dagger})-\sum_{m=0}^{n-2}\tbinom{n-1}{m}\ln(m+\omega_{0})(n-1-m)F(v^{\dagger})^{m}\left(1-F(v^{\dagger})\right)^{n-2-m}f(v^{\dagger}),$
@@
$\ln(m+\omega_{0})-2\ln(m-1+\omega_{0})+\ln(m-2+\omega_{0})=\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)-\left(\ln(m-1+\omega_{0})-\ln(m-2+\omega_{0})\right)<0;$
@@ G-B Step 2: Seller’s Uniform Pricing Analysis
To constitute the PBE, users’ social activity decisions should be consistent with the seller’s belief that is updated according to Bayes’ theorem (i.e., belief consistency [fudenberg1991perfect]). Moreover, anticipating the seller’s uniform price $p_{0}$ in Stage II, users must have no incentive to deviate from the equilibrium strategies of $x_{i}^{*}=1,\forall i\in\mathcal{N}$. This fact then helps guide our proof in this subsection. Rewrite the final payoff for each user $i$ in (4) here:
@@ H-A Impact of $\delta$ on $v^{*}$
    and $\partial^{2}\Theta(v^{*},\delta)/\partial\delta^{2}=2v^{*}>0$, which suggests that $\partial\Theta(v^{*},\delta)/\partial\delta$ increases with $\delta$ on $(2\ln(n-1+\omega_{0})/\bar{v},1)$. We also calculate the following second-order partial derivative:
@@ I-B Proof for Proposition 5
$\sum_{m=1}^{n-1}\sum_{k=1}^{n-1}\frac{(n-2)!}{(n-1-m)!(m-1)!}\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)\leq\sum_{m=1}^{n-1}\sum_{k=1}^{n-1}\frac{(n-2)!}{(n-1-k)!(k-1)!}\left(\ln(m+\omega_{0})-\ln(m-1+\omega_{0})\right)$ (A67)
@@
$\tilde{\Pi}_{0}\triangleq\mathbb{E}_{i\sim\mathcal{N}_{0}}\left(\sum_{i\in\mathcal{N}_{0}}p_{0}d_{i}\right)=$ $|\mathcal{N}|\text{Prob}\left[i\in\mathcal{N}_{0}\right]\cdot\mathbb{E}_{v_{i}\sim f(\cdot|i\in\mathcal{N}_{0})}\left(p_{0}\mathbbm{1}(v_{i}\geq p_{0})\right),$
$=$ $|\mathcal{N}|\left(1-\delta\frac{v^{*}}{\bar{v}}\right)p_{0}\left(1-\frac{1-\delta}{\bar{v}-\delta v^{*}}p_{0}\right),$ (A76)
@@
$\left((\delta-1)\tilde{L}^{\prime}-(2-\delta)\tilde{J}\right)\frac{\bar{v}}{\delta}+4\frac{1-\delta}{\delta^{2}}\tilde{J}\tilde{L}^{\prime}\leq$ $\left(\left((\delta-1)\tilde{L}^{\prime}-(2-\delta)\tilde{J}\right)\frac{\bar{v}}{\delta}+4\frac{1-\delta}{\delta^{2}}\tilde{J}\tilde{L}^{\prime}\right)|_{\bar{v}=2\frac{\tilde{J}}{\delta}},$
$=$ $\frac{2\tilde{J}\tilde{L}^{\prime}}{\delta^{2}}\left((1-\delta)-(2-\delta)\frac{\tilde{J}}{\tilde{L}^{\prime}}\right)<0$ (A93)
@@
In the extended four-stage model, the seller sets the uniform price $p_{0}^{e}$ to maximize the total expected revenue $\tilde{\Pi}$ in (A77), i.e., $p_{0}^{e}=\arg\max_{p_{0}\in[0,\bar{v}]}\tilde{\Pi}(p_{0},v^{*}(p_{0}))$. Also, notice that $p_{0}^{e}$ is uniquely determined (see Appendix K-B3). It then follows that $\tilde{\Pi}^{e}\geq\tilde{\Pi}^{*}$ since $p_{0}^{e}\geq p_{0}^{*}$. Hereafter, we use the superscript $e$ to refer to the extended four-stage model and * for the original three-stage model. Then we turn to the expected profiled revenue $\tilde{\Pi}_{1}$ in (A75), and we will have $\tilde{\Pi}_{1}^{e}\geq\tilde{\Pi}_{1}^{*}$ since $v^{e}\geq v^{*}$. For the expected non-profiled revenue $\tilde{\Pi}_{0}$ in (K-B), we have $\tilde{\Pi}_{0}^{e}=\tilde{\Pi}_{0}^{*}$ due to $v^{e}=v^{*}$ and $p_{0}^{e}=p_{0}^{*}$ if $\bar{v}/2\leq 1/\delta\cdot\ln\left(n-1+\omega_{0}\right)$ (Case I of Proposition 8). Otherwise, we have $\partial^{3}\tilde{\Pi}_{0}(v)/\partial v^{3}=\partial^{3}\tilde{\Pi}(v)/\partial v^{3}<0$ according to Lemma A12. Thus, $\partial\tilde{\Pi}_{0}(v)/\partial v$ is concave in $v$ over $v^{o},\bar{v})$ and $\partial\tilde{\Pi}_{0}(v)/\partial v=\partial\tilde{\Pi}(v)/\partial v-\partial\tilde{\Pi}_{1}(v)/\partial v=\partial\tilde{\Pi}(v)/\partial v-|\mathcal{N}|\delta v/\bar{v}$. Moreover, we have that at $v=v^{o}$:
@@
In this appendix, we evaluate the performance of our proposed pricing mechanism for the seller, through which we will understand the power of harnessing social network benefits (see Section VI-B). To evaluate the performance, we introduce two additional benchmarks adapted from the existing literature (e.g., [conitzer2012hide, 2017Is, zhang2011perils, 2020Voluntary]). Specifically, we outline the two benchmarks for the pricing mechanism along with the one we proposed below.
  • Profile-independent Pricing (PIP) (e.g., [conitzer2012hide, 2017Is, zhang2011perils]): The first benchmark involves the seller’s optimal pricing without differentiation based on user profiles. Variations of this benchmark are commonly adopted in the literature on behavior-based price discrimination. In this case, the seller charges a uniform price of $\bar{v}/2$ to all users.
  • Traditional-profiling-based Pricing (TPP) (e.g., [2017Is, 2020Voluntary]): Another benchmark is wherein the seller personalizes price offerings through traditional user profiling. In this case, users voluntarily disclose private information without considering any social network benefits. Detailed equilibrium analyses are provided in Proposition 7, also named no-social-network benchmark there.
@@
In this appendix, we apply our proposed pricing mechanism to real-world consumer profiles, utilizing an annually-updated consumer behavior and shopping habits dataset from [consumer]. Recall that most of our previous analyses have assumed a uniform distribution for user valuations. However, as we will demonstrate later, this empirical experiment using more general distributions further confirms the robustness of our previous pricing methods and major insights.
@@
Figure A3: The user valuation distribution, the fraction of socially active users and the seller’s optimal uniform price $p_{0}^{*}$ in the PBE (whenever it exists), and the seller’s empirical revenue based on a real-world dataset from [consumer].
@@
The dataset from [consumer] offers consumer information such as demographic details, purchased items, purchase amounts, and previous purchase history. Our analysis focuses specifically on handbag purchases without discounts applied and the corresponding $92$ consumer profiles. [^10] We filter the dataset using the criteria ”Item Purchased = Handbag” AND ”Discount Applied = No.” In particular, we focus on purchases made without any promotional discounts applied. This allows us to exclude the potential psychological effects of price discounts on consumer perceptions and purchase intentions [buyukdaug2020effect, simonson2004anchoring]. Moreover, we associate each consumer’s purchase amount with his product valuation, representing the monetary value of his transaction on the purchased item denoted in United States Dollars (USD). To incorporate the social network benefit in our modeling framework, we further hypothetically consider these $92$ consumers as social network users within a community.